#opnfv-meeting: BGS daily synch

Meeting started by frankbrockners at 16:05:06 UTC (full logs).

Meeting summary

    1. Frank Brockners (frankbrockners, 16:05:14)
    2. Peter Bandzi (pbandzi, 16:05:33)
    3. Fatih Degirmenci (fdegir, 16:05:43)
    4. Tim Rozet (trozet, 16:05:44)
    5. daniel smith (lmcdasm, 16:06:42)
    6. update on deploy.sh: virtual is *almost* working, nodes are deployed as vagrant VMs, one issue with nova-compute that I'm trying to fix (trozet, 16:11:07)
    7. update on deploy.sh: pbandzi has tested and figured out proxy support which will be added (trozet, 16:11:37)
    8. Morgan Richomme (morgan_orange, 16:12:16)
    9. update on LFPOD2: jenkins deployed again with success. morgan_orange re-ran part of the test suite which triggered the ODL bugs (trozet, 16:12:16)
    10. ODL issue not seen with only the tempest smoke test suite. pbandzi to try ODL test suite to see if that triggers the bug (trozet, 16:12:38)
    11. trozet to submit install guide to gerrit today (trozet, 16:13:03)
    12. Discussion on the release readiness call earlier today: Focus would be to document which tests pass and which tests fail (frankbrockners, 16:17:23)
    13. https://wiki.opnfv.org/list_of_tempest_smoke_tests (morgan_orange, 16:17:56)
    14. For those tests which fail, let's understand and document the root cause (i.e. test case wrong, setup wrong, component has an issue) (frankbrockners, 16:18:00)
    15. on POD2 SStill 30 errors but seem linked to 5 errors at the end, some look trivial (ambiguous network => tests done on non fresh system then several network starting with the same name test-* (morgan_orange, 16:18:47)
    16. We have common understanding that we won't be able to fix component issues for now. So focus would be on documenting the issues (e.g. if there are issues with Neutron-ODL interaction - let's just explain what those are) (frankbrockners, 16:18:56)
    17. one other thing related to BGS, but not required for Arno. I saw some commits for the hardware adapter stuff the Fuel guys have been working on. If we can make that common, that would be great. (trozet, 16:21:20)
    18. https://wiki.opnfv.org/opnfv_functional_testing (morgan_orange, 16:22:53)
    19. I saw mail that Setfan is proceeding with installation (pbandzi, 16:30:30)
    20. we had issue with vlan0 but hav workaround now (pbandzi, 16:30:44)
    21. and issue wit public network - foreman set dhcp server on it disturb fuel checks (pbandzi, 16:31:25)

Meeting ended at 16:35:17 UTC (full logs).

Action items

  1. (none)

People present (lines said)

  1. frankbrockners (33)
  2. morgan_orange (24)
  3. trozet (15)
  4. fdegir (12)
  5. pbandzi (7)
  6. collabot (3)
  7. ChrisPriceAB (1)
  8. lmcdasm (1)

Generated by MeetBot 0.1.4.